# WaterPark: A Robustness Assessment of Language Model Watermarking

### v2

@@ G.1 Multi-bit Watermarking
While our study focuses on one-bit watermarkers, for completeness, we also evaluate CTWL (Wang et al., 2024), a multi-bit watermarker. In contrast of one-bit watermarkers that encode only a single bit of information (i.e., whether a given text is watermarked), a multi-bit watermarker can encode multiple bits of information into the watermarked text, such as the generating model, the date of generation, and other details. However, despite its larger information capacity, we find that a multi-bit watermarker is typically less robust compared to one-bit watermarkers, as illustrated in Figure 14 and  15.
Figure 14: TPRs of CTWL (with FPRs fixed as 0.01) on different LLMs (OPT and Llama2) and datasets (C4 and HC3).
@@
The test results in the basic encoding and detection scenario highlight the sensitivity of CTWL to different language models. Although it performs well with the OPT model across both C4 and HC3 datasets, its TPR on Llama2 is extremely low, and it becomes completely ineffective when tested on Llama2 using the HC3 dataset. This is due to the lower model perplexity of Llama2. We conduct additional experiments to compare the model perplexity on the same WikiText dataset, and the results show that Llama2 has a perplexity about 6.15, while OPT has about 12.43. Lower model perplexity results to larger fluctuations in the logits produced by the model, makes it more challenging for watermark injection (e.g., increase smaller logits to exceed larger ones).
@@
G.2 Robustness of Watermarkers on Llama2
Our evaluation in §4 mainly uses OPT-1.3B as the underlying LLM. Here, we present the results on Llama2-7B, a more capable LLM. The results can be found at Table  7
Water marker CLEAN Linguistic Variation Textual Integrity Content Manipulation Paraphrase
@@ G.2 Robustness of Watermarkers on Llama2
Table 7: Resilience of LLM watermarkers against attacks on Llama2.
G.2.1 Linguistic Variation Attack
Comparing with the OPT results, most of the watermarkers show TPR drops on Llama2. One exception is GO, which still maintains resilience across all three attacks.
G.2.2 Lexical Editing Attack
We have the following observation in Llama2: i) Context-free watermarkers (e.g., UG) show significantly higher attack resilience against swapping attacks. This may be due to the fact that the attack alters the word and sentence order, which affects the detection of both text-dependent and index-dependent watermarkers. ii) Similar to OPT, text-dependent, (hard) distribution-reweighting watermarkers tend to be more vulnerable to lexical editing attacks. iii) The typoing and swapping attacks show significantly higher attack effectiveness on Llama2, similar to our observations on OPT.
G.2.3 Text-mixing Attacks
With the exception of watermarkers based on distribution transform (e.g., RDF and GO), which shows greater resilience to text-mixing attacks, the other methods are virtually undetectable, which is consistent with the observations in OPT
G.2.4 Paraphrasing Attack
All methods show a significant TPR decrease, only GO presents a remarkable resilience. RDF performs well in OPT and does not present a significant resilience on Llama2.
G.3 Fidelity Preservation of attacks
@@ Appendix H Additional Related Work
LLM watermarking In response, LLM watermarking emerges as a promising alternative, which instruments the LLM generative process with statistical signals that can be subsequently detected. The existing LLM watermarking techniques can be categorized based on the stages in which they are appliedLiu et al. (2023b): i) training-time watermarkingLiu et al. (2023c); Tang et al. (2023); Sun et al. (2022, 2023); Xu et al. (2024), ii) watermarking during logit-generationKirchenbauer et al. (2023a); Zhao et al. (2023); Hu et al. (2024); Liu et al. (2023a, 2024); Fairoze et al. (2023); Ren et al. (2023); Fernandez et al. (2023); Wang et al. (2024); Yoo et al. (2023); Lee et al. (2023); Giboulot and Teddy (2024), iii) watermarking during token-sampling Kuditipudi et al. (2023); Aaronson and Kirchner ; Hou et al. (2023), and iv) post-generation watermarkingZhang et al. (2023); Yoo et al. (2024); Yang et al. (2023); Munyer and Zhong (2023); Sato et al. (2023); Abdelnabi and Fritz (2021). This study mainly focuses on training-free, pre-generation watermarking, which applies to any given LLMs and provides flexible control over multiple criteria (e.g., quality, effectiveness, and robustness).
@@
Evaluation of LLM watermarkers. As LLM watermarkers become prevalent, recent work attempts to benchmark the performance of various watermarkers. However, current studies either focus solely on the effectiveness of watermarking or have limited assessments of robustness. For instance, WaterBenchTu et al. (2023) compares the effectiveness of TGRLKirchenbauer et al. (2023a) and UGZhao et al. (2023) under varying hyper-parameter settings (e.g., prompt length); LLM-judgerSingh and Zou (2023) uses GPT-3.5-Turbo as a judge to evaluate the effectiveness of RDFKuditipudi et al. (2023) and TGRLKirchenbauer et al. (2023a) and employs a binary classifier based on MLP to distinguish between watermarked and non-watermarked texts; MarkMyWordsPiet et al. (2023) compares the effectiveness of four watermarkers including TGRLKirchenbauer et al. (2023a), GOAaronson and Kirchner , RDFKuditipudi et al. (2023), and UWChrist et al. (2023) and only evaluates the robustness of TGRL against watermark removal attacks.
To our best knowledge, this is the first study dedicated to the robustness of LLM watermarkers against watermark removal attacks. We aim to understand how different design choices impact watermarkers’ attack robustness and to identify best practices for operating watermarkers in adversarial environments.

### v3

@@ G.1 Multi-bit Watermarking
While our study focuses on one-bit watermarkers, for completeness, we also evaluate CTWL (Wang et al., 2024), a multi-bit watermarker. In contrast of one-bit watermarkers that encode only a single bit of information (i.e., whether a given text is watermarked), a multi-bit watermarker can encode multiple bits of information into the watermarked text, such as the generating model, the date of generation, and other details. However, despite its larger information capacity, we find that a multi-bit watermarker is typically less robust compared to one-bit watermarkers, as illustrated in Figure 12 and  13.
Figure 12: TPRs of CTWL (with FPRs fixed as 0.01) on different LLMs (OPT and Llama3) and datasets (C4 and HC3).
@@
The test results in the basic encoding and detection scenario highlight the sensitivity of CTWL to different language models. Although it performs well with the OPT model across both C4 and HC3 datasets, its TPR on Llama3 is extremely low, and it becomes completely ineffective when tested on Llama3 using the HC3 dataset. This is due to the lower model perplexity of Llama3. We conduct additional experiments to compare the model perplexity on the same WikiText dataset, and the results show that Llama3 has a perplexity about 6.15, while OPT has about 12.43. Lower model perplexity results to larger fluctuations in the logits produced by the model, makes it more challenging for watermark injection (e.g., increase smaller logits to exceed larger ones).
@@
G.2 Additional Robustness of Watermarkers
Here we present the full robustness evaluation results across different models and datasets, which strongly corroborate our primary experimental findings. These experiments revealed several significant patterns that align with earlier observations, demonstrating that the relative robustness rankings of watermarking methods remain consistent regardless of model architecture or dataset characteristics. This consistency suggests that robustness is primarily determined by the inherent algorithmic properties of each method.
Table  9 shows the detailed results. Notably, UPV and SIR exhibited substantial vulnerability, with watermark detectability degrading even under simple Linguistic variation and Lexical editing attacks, indicating fundamental limitations in their robustness across model scales. Text-mixing attacks posed significant challenges for most watermarkers—particularly UPV, which failed completely under such conditions—while RDF and GO maintained reliable detection, highlighting key differences in algorithmic resilience. Furthermore, under paraphrasing attacks, both UG and RDF consistently achieved high detection rates, further supporting the conclusion that inherent watermarking design, rather than model- or dataset-specific factors, governs robustness against adversarial manipulations.
@@ G.2 Additional Robustness of Watermarkers
QWEN + Paper Conclusion
TGRL 0.993 0.944 0.944 0.831 0.429 0.222 0.887 0.915 0.000 0.667 0.833 0.833 0.667 0.485 0.222
UG 0.993 0.857 0.833 0.976 0.929 0.930 0.976 0.738 0.857 0.714 0.714 0.991 0.976 0.877 0.921
UPV 0.400 0.400 0.400 0.430 0.080 0.000 0.380 0.360 0.000 0.000 0.000 0.000 0.200 0.080 0.240
RDF 0.999 0.998 0.996 0.996 0.976 0.979 0.991 0.993 0.872 0.893 0.932 0.978 0.905 0.738 0.978
UB 0.980 1.000 1.000 0.962 0.000 0.033 0.921 1.000 0.018 0.042 0.000 0.485 0.514 0.103 0.255
SIR 0.978 0.580 0.500 0.420 0.320 0.320 0.440 0.560 0.340 0.360 0.360 0.460 0.400 0.300 0.000
GO 0.996 0.996 0.996 0.994 0.982 0.956 0.986 0.996 0.864 0.887 0.946 0.982 0.640 0.560 0.667
G.3 Different Styles Datasets
To comprehensively evaluate watermarking performance across different text generation scenarios, we created diverse datasets spanning multiple styles and domains. The Paper Conclusion dataset was constructed by collecting research papers from arXiv and having LLMs summarize their key findings. We also incorporated Law Stack Exchange, a legal domain question-answering dataset sourced from the Stack Exchange forum, and WritingPrompts, a creative writing dataset focused on story completion tasks. These datasets can be categorized into three distinct styles: completion-style (C4 and Story Completion datasetsEuclaise ), question-answering (HC3 and Law Stack Exchange datasetsLi ), and summarization (Paper Conclusion dataset). Our analysis revealed several significant findings. Table  10 shows the results.
First, all watermarkers demonstrated notably lower performance on question-answering style datasets. This degradation was particularly pronounced for UPV and SIR, which achieved only around 0.5 TPR on these datasets. This can be attributed to the prevalence of domain-specific terminology and fixed expressions in Q&A datasets, which inherently constrain the model’s linguistic choices and make watermark embedding more challenging. Second, robust watermarkers like RDF and GO maintained exceptional performance across all dataset styles, demonstrating their versatility and reliability. RDF achieved consistently high TPR scores ranging from 0.934 to 0.999, while GO demonstrated even more stable performance with TPR values between 0.978 and 0.998. These results highlight an important pattern: watermarkers that demonstrate strong baseline performance maintain their effectiveness across different text generation styles, while less stable or lower-performing watermarkers are particularly vulnerable to degradation when applied to question-answering formats. Notably, the high watermarking effectiveness on Paper Conclusion tasks (TPR > 0.99 for most methods) can be attributed to the inherent flexibility in academic summarization. Unlike Q&A tasks that require specific terminology and fixed expressions, paper conclusions allow for more freedom in expression while conveying the same core findings. This flexibility in language choice and sentence construction provides more opportunities for watermark embedding without compromising the semantic accuracy of the summary.
Water marker Summarization Completion Question-Answer
Conclude C4 StoryComple HC3 LawQA
TGRL 0.993 0.992 0.992 0.926 0.934
UG 0.993 0.972 0.981 0.864 0.891
UPV 0.400 0.540 0.560 0.460 0.430
RDF 0.999 0.998 0.996 0.945 0.934
SIR 0.978 0.965 0.976 0.566 0.631
GO 0.996 0.998 0.994 0.986 0.978
Table 10: Performance of watermarkers across different datasets on Qwen2.5-14B.
G.4 Fidelity Preservation of attacks
@@ Appendix H Additional Related Work
LLM watermarking In response, LLM watermarking emerges as a promising alternative, which instruments the LLM generative process with statistical signals that can be subsequently detected. The existing LLM watermarking techniques can be categorized based on the stages in which they are appliedLiu et al. (2023b): i) training-time watermarkingLiu et al. (2023c); Tang et al. (2023); Sun et al. (2022, 2023); Xu et al. (2024), ii) watermarking during logit-generationKirchenbauer et al. (2023a); Zhao et al. (2023); Hu et al. (2024); Liu et al. (2023a, 2024); Fairoze et al. (2023); Ren et al. (2023); Fernandez et al. (2023); Wang et al. (2024); Yoo et al. (2023); Lee et al. (2023); Giboulot and Teddy (2024), iii) watermarking during token-sampling Kuditipudi et al. (2023); Aaronson and Kirchner ; Hou et al. (2023), and iv) post-generation watermarkingZhang et al. (2023); Yoo et al. (2024); Yang et al. (2023); Munyer and Zhong (2023); Sato et al. (2023); Abdelnabi and Fritz (2021). This study mainly focuses on training-free, pre-generation watermarking, which applies to any given LLMs and provides flexible control over multiple criteria (e.g., quality, effectiveness, and robustness). Some watermarks also consider asymmetric schemes (Fairoze et al., 2023), while our study mainly focuses on symmetric watermarking schemes due to their widespread adoption.
@@
Evaluation of LLM watermarkers. As LLM watermarkers become prevalent, recent work attempts to benchmark the performance of various watermarkers. However, current studies either focus solely on the effectiveness of watermarking or have limited assessments of robustness. For instance, WaterBenchTu et al. (2023) compares the effectiveness of TGRLKirchenbauer et al. (2023a) and UGZhao et al. (2023) under varying hyper-parameter settings (e.g., prompt length); LLM-judgerSingh and Zou (2023) uses GPT-3.5-Turbo as a judge to evaluate the effectiveness of RDFKuditipudi et al. (2023) and TGRLKirchenbauer et al. (2023a) and employs a binary classifier based on MLP to distinguish between watermarked and non-watermarked texts; MarkMyWordsPiet et al. (2023) compares the effectiveness of four watermarkers including TGRLKirchenbauer et al. (2023a), GOAaronson and Kirchner , RDFKuditipudi et al. (2023), and UWChrist et al. (2023), and evaluates the robustness of TGRL against watermark removal attacks.
While our study shares some evaluation components with MarkMyWords, our contributions differ substantially in scope and depth. MarkMyWords includes robustness testing as part of a broader evaluation, whereas our work provides a dedicated and systematic framework specifically for stress-testing watermark robustness. This specialized focus enables deeper exploration of attack types, adaptive strategies, and evaluation metrics. Methodologically, we introduce several new dimensions not covered in prior work, such as causal analysis of watermark design choices, comparisons between specific and generic detectors, and adaptive attacks mounted by adversaries using generic detectors. Furthermore, we establish a standardized benchmark comprising 12 representative attacks and 8 evaluation metrics—facilitating fine-grained robustness analysis, including next-token distribution comparisons and what-if attack simulations.
Our systematic approach has uncovered several novel findings that go beyond the scope of existing evaluations—for example, challenging assumptions in prior work (see Table 1) and identifying key trade-offs in design robustness. We acknowledge the need to better articulate these distinctions and will revise the related work section to more clearly position our contributions relative to prior efforts such as MarkMyWords.
To our best knowledge, this is the first study solely dedicated to evaluating the robustness of LLM watermarkers against removal attacks. Our goal is to understand how design decisions influence resistance to adversarial perturbations and to establish best practices for deploying watermarkers in real-world adversarial environments.
@@ Appendix I Evaluation Guidelines for Future LLM Watermarking Research
Appendix J Deployment Guidelines for Adversarial Environments
To complement our empirical evaluation, we distill our findings into actionable deployment guidelines for practitioners operating watermarkers in adversarial settings.
Watermarker selection. Our results show that watermarkers differ significantly in their robustness to specific attack types. We recommend selecting watermarkers based on the anticipated threat landscape: RDF and UG are especially resilient against paraphrasing and text-mixing, while TGRL performs well against linguistic variations. Practitioners should avoid UPV and SIR in high-adversity settings due to their instability and vulnerability under attack combinations.
Detector configuration. For watermarkers with unstable detectors (e.g., UPV), using hybrid detection schemes—combining score-based and model-based detectors or incorporating a generic detector—can mitigate failure modes. Our results (Table  3) suggest that generic detectors can complement specific ones, enhancing recall without significant degradation in precision.
Multi-layered defense. Since individual attacks (e.g., typoing) are often insufficient to remove watermarks, while combined or adaptive attacks (e.g., swapping + synonymizing or GBDA) are far more effective, we recommend adopting multi-layered detection pipelines. These may involve edit-distance-aware scoring (e.g., RDF’s edit score) and robustness testing against common perturbation chains.
Risk assessment. To assess real-world risk, practitioners should consider not only overall TPR/AUC, but also worst-case TPR under fixed low FPR (e.g., 1%) across combined and high-intensity attacks. Our findings highlight that even robust watermarkers may fail under iterated paraphrasing (Figure  7).
Resource-aware deployment. For cost-sensitive applications (e.g., online moderation), we recommend avoiding watermarkers requiring expensive generation/detection steps (e.g., deep model-assisted detection in UPV), and instead favoring methods like UG or RDF that are both robust and lightweight.